# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ III.2.1 Positive interactions and perceptions
Chloe explicitly says that her interactions with her peers have been good and that she does not think that her gender excludes her in spaces: “I’ve never really felt excluded because of my gender and that I’ve always been in classes with other girls and-but I will say I have made groups for all my study groups and people I really talk to are other girls and that has been the most comfortable for me.” It may be that Chloe has surrounded herself with supportive people, specifically women who tend to be more approachable, thus avoiding any negative interactions and experiences with male peers. For instance, her study groups consist of other women.
@@
Ruby is the only one who explicitly mentions that men and women work together well. In particular, she mentions not worrying about asking male peers for help and how she perceives her interactions with them to be positive: “I think [my college] has a very good environment in terms of men and women working together and collaborating. I have, I really haven’t felt that fear of asking a guy for help versus asking a girl for help here as much as in high school.” Despite these positive reflections, we note that she is speaking about her college experiences relative to her high school interactions. In the next section, we will discuss negative experiences Ruby had in college with male peers.
III.2.2 Negative interactions and perceptions
@@
All of the women recalled negative experiences with their male peers during college. At the end of this section, we include Paulina’s experiences as separate from the others, but still under this theme. The reason for doing so is that her experiences are different due to her specifically pointing to both her race and gender as being aspects of her identity that led to her negative experiences with both male and female students.
@@
Cathy did not provide many experiences about her peers in physics but describes that she has had many negative experiences with peers in her computer science courses. She says: “I think part of the problem is that, in the computer science classes, a lot of people come in with a lot of experience because there’s so many side projects you can do on your own and learn stuff on your own, so, and everybody has to take the intro classes you can’t leave-place out, so I went into my first intro class, not knowing any computer science and there were people [who] would just do the lab in 10 seconds and then leave because they’ve been doing it for years and they just had to fill the requirement. So I think that made it a little more intimidating because you would ask someone for help, but they’d be already out the door and they didn’t really have time to slow down”. According to Cathy, peers in her computer science courses are intimidating to approach to ask for help since it seems like they have a lot of background in computer programming which is very relevant in the courses. Without any actions from the instructor to put students in groups in which they can learn from each other on projects, the structure of these courses did not support peer collaboration, which made it more difficult to connect with other students.
@@
Chloe shares an experience in which she had an exam and following the exam, she heard some students talk negatively about it: “a bunch of students were in the hallway and then they’re all…talking about how they didn’t like the exam and they didn’t like the format of the exam and that how…they just didn’t like it, and it was kind of, I felt bad vibes from the conversation, and most of them were boys.” Chloe’s bad vibes about male students complaining about the exam reflects how bad she felt about them not respecting the type of exam the instructor had given and their excessive complaining did not create a welcoming environment for Chloe in which she would feel comfortable communicating with them. This may also illustrate how men may be more critical of assessments including their format, especially if it was given by a female instructor.
@@
In regard to working in groups, Chloe says that she does not feel as comfortable working with men as she does working with women: “I just always felt a little more uncomfortable [with male peers] and I’m not sure how much of that well, I guess it’s all just an inherent thing that’s still part of our culture… I know that’s a difficult problem to solve.” According to Chloe, she mostly interacts with female peers within study groups and feels comfortable doing so. When it comes to working with male peers, she says, “I’d only really feel comfortable working with the boys if I’m already friends with them outside of the classroom, like they’re one of my closer friends.” It seems like for men, there’s a higher threshold to overcome in order for them to be deemed comfortable enough for Chloe to work with them.
Chloe also shares that although her physics department is not overly dominated by men, “it’s all like males at the top.” She adds: “I feel like as a woman, I have to prove myself for people to look at me and go, Oh, you’re smart and dedicated and you like science, I have to prove that first. It’s not implicitly assumed and nobody assumes that when they first meet me. Well, I think, for many men in my classes, it’s the opposite where it’s like you initially assume that this person is really smart and gets really good grades just because they’re male in science.” Thus, Chloe points to a pervasive gender stereotype in physics that for men, everyone assumes that they are smart and will do well in physics (or other STEM disciplines). On the other hand, for women, this is not assumed and she has to always prove this to people first.
@@
When asked about who is part of her study group, Francis explains that she mostly works on homework with female students because of the negative experiences she’s had with male peers: “I would generally choose a woman just because…I remember that last semester we would, in one of my physics classes, if me and my partner had different answers, it kind of felt like he’d automatically assume that he was right, which wasn’t always the case, and to me that just felt like a very male thing to [do]- So I guess I do generally would select to study with a woman.” Thus, having negative experiences with men in her study groups has affected whom she currently chooses to work with. She also attributes this behavior of automatically assuming they have the correct answer as a “very male thing to do”, which hints to a masculine physics culture.
@@
She also says that men have more confidence in answering questions and their confidence takes up space in the classroom: “it felt like there was a lot more men in the class…than [there actually] are…most of the speaking seem to be coming from them and I guess, they just felt more confident answering questions whereas I think, for me, I wouldn’t answer a question unless I know for sure that I’m right, but they’re happy just to have a go with something that they think could be right.” This correlates with prior research in which men in physics courses are more confident in answering questions [23, 36]. When asked if she thinks that these students could be showing off, she says, “I definitely do get that idea from some people. I remember one guy who would just, he would ask questions that kind of felt like he knew the answers [to]. He just wanted to show that he was thinking about this one specific thing.” This type of report of male students showing off reflects the physics culture in which men feel comfortable doing so. Regardless of whether this was intentional or not, women tend to perceive this behavior as intimidating or these types of behaviors can reduce their sense of belonging.
Karina shares that she has a different approach to learning, she takes longer to learn the material. She also explains how this negatively affects her interactions with her male peers: “I need to think about things for a little bit to get a really good understanding of them and that’s how I work, so I think probably in labs I sometimes felt like the group was moving on or they had done the thing and I didn’t want to speak up…maybe just because of pride or something. Also, I didn’t want to be like oh, hold on, I don’t understand what’s going on here, can I try this, can I do this, and I’m not sure how much of that has to do with gender versus, you know, just the type of learner I am and the way I work.” Thus, while on the one hand, Karina seems to be contemplating as to whether her gender was a reason for these negative interactions and the fact that she did not feel comfortable speaking up about letting her try certain things and slowing down, she also puts the onus on herself and contemplates whether it is her learning style that is making it difficult for her to collaborate with her male peers in the physics labs and giving her inadequate opportunities to delve deeper into the concepts and be on the same page with them.
@@
When asked about what the vibes were like in the classroom and if there was a dominant group, Karina reflects: “more participation, more dominant, men are, they’re…sort of like blocking out the space to ask more basic questions to understand the material at the most basic level.” She further provides a similar insight as Francis did: “I think there’re some spaces, where I don’t know if this is a gender thing, but there would be men asking like questions that perhaps are built to make himself seem really smart and so it’s suffocating this space to ask sort of more basic level questions for other people.” Again, Karina adds to Francis’s point of men often showing off and wanting to be perceived as smart by asking certain type of questions, by saying that they are “suffocating the space” for other people to ask more basic questions. From her phrasing, we interpreted these types of behaviors to be very negative experiences. These behaviors also alienate and make other students anxious by creating overly high expectations for types of questions that can be asked or knowledge that can be brought to the table during a physics class.
@@
Although Ruby shared previously that she perceives the collaborative environment to be positive, she also shares that her interactions with her male peers are varied: “A lot of the guys in my class are so helpful and so willing to help and don’t approach things in condescending manner at all, but then there are others who act like know it all…I don’t know if they do it because I’m a woman, but sometimes it feels like that, where they’re like mansplaining so often in and outside the classroom, there is definitely a culture of mansplaining.” For example, she recalls an experience working with a male peer lab partner via Zoom. She says that her partner was extremely condescending and, “I don’t know if it was because I was girl, I think he just thought he was smarter than everyone else in the room at all times and always had to be the one to answer the question first, always had to be the one to know it…I ended up switching my partners actually because I just could not deal with that.”
@@
While on Zoom, she says that her partner would dominate breakout room conversations and assignments, leaving her and her other female partner behind: “He would always just solve it himself and then present the answer, without even giving us a chance to try and that really bothered me because it made me feel like…he thinks oh, because he knows it, that all of us cannot be smart enough to even try…and it made me feel like I was being talked down to from him, thinking that I was less smart than he was.” The fact that this happened over Zoom may have made this situation especially difficult to be in. Ruby also adds that her partner would not collaborate on the written report and would write it himself. She says, “he would act like I didn’t know anything that I was doing, when I was fully capable of doing it myself and that really made me feel…angry about it because… I love collaboration, I think it’s so important, and I really liked when I’m able to work with somebody, we’re able to combine our own knowledge and figure out new stuff, and with him, I felt that I was limiting my own knowledge, because I wasn’t getting access to even try. And I felt he was mansplaining things all the freaking time. I really did not like it.”
@@
When asked if she thought that this behavior was specifically towards women, Ruby says, “100% yes.” Again, she explains that she thought that the student thought that he was smarter than others, “but he also believed that he was helping in a way. He thought that, Oh, these women can’t do it themselves, I’ll help them, I’ll do it for them, which is so not how real life works and that’s not how we should handle the situation at all. But in my mind at least, I believe that’s why he did, that I think he thought of women in a sense, that they were less capable and more vulnerable than men, and so he wanted to be this savior.” Thus, in her experience, this male peer thought that he was helping women by answering for them and doing all the work, when in reality, it had the opposite effect.
@@
Cathy recalls her physics REU experiences at large research universities, some of them being very negative. In one of her REUs, she was the only woman in a group of nine students and she was not familiar with the research topic. As a result, she felt excluded: “I couldn’t really ask questions or anything because everybody else knew so much more than me and they kind of talked down to me, sometimes it was just really overwhelming and intimidating and it definitely harmed my confidence a lot, because since I didn’t feel like I could ask questions, I really didn’t understand what was going on, and I got super behind, which made it so it was harder to ask questions, it’s just kind of the cycle and I don’t really feel like I learned anything in the end because I was just so overwhelmed the whole time.” Without any support from REU mentors, the fact that the other male students gave her the impression that they were more knowledgeable than her and they talked down to her made her intimidated to ask questions, potentially making her not benefit from the entire REU experience when a supporting environment would have made it very beneficial.
III.2.3 Paulina as an island among her college peers
Chloe recalls during high school when her history teacher made an observation about the physical distribution and clustering of students in the classroom: “We made a very big point in class about the us versus them…ideologies and they pointed out how in our classroom, every single table was separated by gender and race…and that we all gravitated towards people who were similar, who looked similar to us and that it was so clear…the seating arrangements in the classroom…” This excerpt is also relevant because Paulina emphasized that she was particularly excluded in her classes due to the intersection of both her gender and race as described below.
@@ III.2.3 Paulina as an island among her college peers
She says that being a woman of color is complicated because, “whenever I think about asking someone for help…it’s like well, people perceive me as being different in the classroom because I’m a person of color and also because I’m a woman. So, there’s these two factors that they can visually see when I approach them and I always feel like that’s…the first thing they see and it’s how they assess my value or what I can contribute to the classroom.” As one of the few people of color in her program, we note that this may be extremely difficult to navigate as she constantly thinks about how others are perceiving her. She also says that students simultaneously are judging her and assessing her intellect and the value of interacting with her and what she will bring to collaboration of physics problem solving. Similar issues have been documented in previous studies involving interviews with six physics graduate students who were Black women [21].
Paulina explains that she even feels judged amongst other women in her course, “even amongst a group of other women, there’s also been this expectation that I am just not up to par with them, because [of] my racial background and so I feel there’s some intersections in terms of my experiences in STEM with women…I definitely know that beyond the realm of just gender, race also plays a really big role in it.” In Paulina’s experience, she doesn’t necessarily find solidarity amongst other White women either because she is a woman of color. Since her university is both small and a PWI, finding support and solidarity in other students who look like her can be extremely difficult.
@@
Being ignored by peers: Much of Paulina’s experiences are centered around her working alone on assignments because she is ignored by her peers during group work. Below, we discuss how her intersecting marginalized identities negatively impact her throughout and play a central role in her interactions with peers. Paulina explains that her peers will not value or listen to her contribution in groups: “A lot of my experiences in the classroom have been where people, often peers, for example, will kind of look at me and if we have to engage in sort of group work, there’s usually not a lot of listening happening on their behalf. I think there’s this assumption that… because I am a woman of color in STEM, I don’t necessarily have the same background or knowledge as they [do], so a lot of the times I’ll bring [up] points and it’ll just be ignored and that’s something that I’ve had to deal with a lot.” Here we see that her peers are ignoring her, but she recognizes and adds that this could be due to biases that they have about her physics background based on her being a woman of color. Thus, stereotypes and biases impact how other students perceive her, how she perceives her peers perceiving her, and ultimately why other students in physics choose not to engage with her.
She reveals that this semester for the first time, she has been placed in a group with students who she has worked with before in her class. These students have similar physics backgrounds to Paulina and she feels that although there seems to be some effort to meet and work through assignments, it is difficult. She reflects upon working with these students outside of class that she has interacted with in class earlier, “…it’s made me feel more comfortable but usually it just ends up happening, where there’s a lot of physics majors and those that are physics majors are like an established group or they’re just not really open to working with new people outside of the classroom, so it becomes difficult to set that up.” Her experiences reveal that physics students form cliques with others with similar identities which makes it difficult for her to work with them because she does not share their identity. She says, “[t]hey either stick to doing things on their own, or like they already have a background with other students in the class, so they’ll just reach out to them, and if we happen to be in the same group, they’ll usually just choose to staying engaged with the students they already know, rather than work together as a group with me who they might not be as familiar with.” Thus, according to Paulina, other students feel more comfortable with others who share their identity and whom they know more personally or have interacted with more. The fact that these students have all been placed in a group with Paulina to work outside of the class and she is still feeling excluded (both inside and outside of the classroom) is very concerning. This formation of a clique makes it difficult for Paulina to meaningfully collaborate with other peers (including other women) as they choose to talk to other students in other groups who they do know and have similar identities (e.g., gender and race), as opposed to her, who is placed in their group.
@@
Paulina explains that although some of her classes incorporate group work, there is no effort to communicate or work on problems outside of class. Even for the case mentioned earlier in which she has been placed in a group for studying outside the class, she feels that it does not work effectively, and she is left to work on her own. She says, “Usually what ends up happening is that we don’t really communicate with each other through the group messaging, for example, and so it ends up being a situation where no one, or at least I’m aware of, no one is reaching out to each other to talk and work things out together so by default I just feel like I have to end up doing things on my own.” Thus, even when she has been placed in a group to work outside of class, her own group does not communicate about when they should get together to study or do homework, so she resorts to working alone. It is unclear whether students who have been placed in Paulina’s group are meeting with other students with similar identity outside of the class to study and do homework. In particular, she is not aware of what is happening since nothing happens through the messaging system they were supposed to communicate through when they were placed in the same group for out of class study. Whether this lack of communication is directed towards Paulina because of her identity or caused by other reasons such as none of the students in the group she has been placed meeting regularly, it is interpreted as lack of community by her. Thus, she is unable to lean on her peers as resources and collaborators, and can only rely on herself.
@@
Inside the classroom, when she is ignored by her peers, she feels like she cannot rely on her group mates either and says, “I felt like I could only directly reach out to the professor to ask questions or I was expected to learn things on my own and if I didn’t verify [that my problem solving process was correct with the professor] it reflected badly on my own personal knowledge and understanding of the subject. And it’s still something that I’m trying to grapple with, but a lot of my experiences usually have to do with people underestimating me or feeling like I don’t have enough like a solid background in physics, and so it causes me to feel like I have to continuously overachieve and prove that I am just as knowledgeable as everyone else in the space.” Thus, Paulina describes that in her case, she has to self-study, rather than work in a group to learn the material or work on problems together. She also adds that she has to understand the material well even though she is not able to collaborate with her peers. She also feels that if she does not “verify” with the instructor that she has learned the material well and solved the associated problems using the correct approach, then it makes her look bad. The fact that she feels that she has to overachieve and prove that she is knowledgeable like her peers is an additional burden on her because she does not want to confirm any stereotype others may have related to her identity [61].
@@
Because other students would not work with her even in class even though she was in a group with them even when the structure of her classes utilized group work, she could not complete in-class assignments during class like her peers did. She says: “I would fall behind because I would basically be left to my own devices to work out these problems and then, because I wouldn’t have the time, the resources to ask questions like my peers, I would have to go back to my room and work on it on my own where like my peers would be able to have a conversation and just finish it in the classroom.” Not only does she have to manage feelings of being excluded and negatively stereotyped by her peers, she also has to spend extra time outside of class to finish her class assignments. We can see how her multiple marginalized identities in physics can take up emotional and mental space that could be used for learning physics and problem solving. She has to use additional time and resources to complete her in-class assignments due to the exclusion by peers and navigate her physics journey alone.
@@
III.3 Interactions with college instructors
In this section, we describe the experiences these women had with their college physics instructors. These experiences include Positive and Negative interactions and perceptions.
III.3.1 Positive interactions and perceptions

### v2

@@ IV.2.1 Positive interactions and perceptions
Chloe explicitly says that her interactions with her peers have been good and that she does not think that her gender excludes her in spaces. She clarifies that she has always been in classes with other girls, so this could have been a possible reason for her feeling included. With regard to forming groups, Chloe explains, “I have made groups for all my study groups and people I really talk to are other girls and that has been the most comfortable for me.” It may be that Chloe has surrounded herself with supportive people, specifically women who tend to be more approachable, thus avoiding any negative interactions and experiences with male peers. For instance, her study groups consist of other women.
@@
Ruby is the only one who explicitly mentions that men and women work together well. In particular, she mentions not worrying about asking male peers for help and how she perceives her interactions with them to be positive: “I think [my college] has a very good environment in terms of men and women working together and collaborating…I really haven’t felt that fear of asking a guy for help versus asking a girl for help here as much as in high school.” Despite these positive reflections, we note that she is speaking about her college experiences relative to her high school interactions. In the next section, we will discuss negative experiences Ruby had in college with male peers.
IV.2.2 Negative interactions and perceptions
@@
All of the women recalled negative experiences with their male peers during college. At the end of this section, we include Paulina’s experiences as separate from the others, but still under this theme. The reason for doing so is that her experiences are different due to her specifically pointing to both her race and gender as being aspects of her identity that led to her negative experiences with both men and women students.
@@
Cathy did not provide many experiences about her peers in physics but describes that she has had many negative experiences with peers in her computer science courses. According to Cathy, peers in her computer science courses are intimidating to approach to ask for help since it seems like they have a lot of background in computer programming which is very relevant in the courses. She also notes that for computer science courses, “a lot of people come in with a lot of experience,” also making it difficult to reach out to peers. Without any actions from the instructor to put students in groups in which they can learn from each other on projects, the structure of these courses did not support peer collaboration, which made it more difficult to connect with other students.
@@
Chloe shares an experience in which she had an exam and following the exam, she heard some students talk negatively about it. For example, they did not like the format of the exam. She says that she “felt bad vibes from the conversation, and most of them were boys.” Chloe’s bad vibes about male students complaining about the exam in a course taught by a woman instructor reflects an uncomfortable environment. This may also illustrate how men may be more critical of assessments including their format, especially if it was given by a woman instructor.
@@
In regard to working in groups, Chloe says that she does not feel as comfortable working with men as she does working with women. She also reflects on why this may be the case: “well, I guess it’s all just an inherent thing that’s still part of our culture… I know that’s a difficult problem to solve.” According to Chloe, she mostly interacts with women peers within study groups and feels comfortable doing so. When it comes to working with male peers, she says, “I’d only really feel comfortable working with the boys if I’m already friends with them outside of the classroom, like they’re one of my closer friends.” It seems like for men, there’s a higher threshold to overcome in order for them to be deemed comfortable enough for Chloe to work with them.
@@
When asked about who is part of her study group, Francis explains that she mostly works on homework with women because of the negative experiences she has had with male peers: “in one of my physics classes, if me and my partner had different answers, it kind of felt like he’d automatically assume that he was right, which wasn’t always the case, and to me that just felt like a very male thing to [do]- So I guess, I generally would select to study with a woman.” Thus, having negative experiences with men in her study groups has affected whom she currently chooses to work with. She also attributes this behavior of automatically assuming they have the correct answer as a “very male thing to do”, which hints to a masculine physics culture.
@@
She also says that men have more confidence in answering questions and their confidence takes up space in the classroom, saying that it felt like “there was a lot more men in the class than there actually are”. In contrast, she does the opposite: “I wouldn’t answer a question unless I know for sure that I’m right…” This correlates with prior research in which men in physics courses are more confident in answering questions [23, 36]. When asked if she thinks that these students could be showing off, she says, “I remember one guy who… would ask questions that kind of felt like he knew the answers [to]. He just wanted to show that he was thinking about this one specific thing.” This type of report of male students showing off reflects the physics culture in which men feel comfortable doing so. Regardless of whether this was intentional or not, women tend to perceive this behavior as intimidating or these types of behaviors can reduce their sense of belonging.
@@
When asked about what the vibes were in the classroom and if there was a dominant group, Karina recalls that men were more dominant and they participated more. As a result, she says, “they’re…sort of like blocking out the space to ask more basic questions to understand the material at the most basic level.” She further provides a similar insight as Francis did: “there would be men asking like questions that perhaps are built to make himself seem really smart and so it’s suffocating this space to ask sort of more basic level questions for other people.” Again, Karina adds to Francis’s point of men often showing off and wanting to be perceived as smart by asking certain type of questions, by saying that they are “suffocating the space” for other people to ask more basic questions. From her phrasing, we interpreted these types of behaviors to be very negative experiences. These behaviors can also alienate and make other students anxious by creating overly high expectations for types of questions that can be asked or knowledge that can be brought to the table during a physics class.
Karina shares that she has a different approach to learning, she takes longer to learn the material. She also explains how this negatively affects her interactions with her male peers: “in labs I sometimes felt like the group was moving on… and I didn’t want to speak up…maybe just because of pride or something. Also, I didn’t want to be like ‘Oh, hold on. I don’t understand what’s going on here.’, ‘Can I try this?’, ‘Can I do this?’ and I’m not sure how much of that has to do with gender versus, you know, just the type of learner I am and the way I work.” While on the one hand, Karina seems to be contemplating as to whether her gender was a reason for these negative interactions and the fact that she did not feel comfortable speaking up about letting her try certain things and slowing down, she also puts the onus on herself and contemplates whether it is her learning style that is making it difficult for her to collaborate with her male peers in the physics labs and giving her inadequate opportunities to delve deeper into the concepts and be on the same page with them.
@@
Although Ruby shared previously that she perceives the collaborative environment to be positive, she also shares that her interactions with her male peers are varied. For example, many of the men in her class, she describes as helpful, “willing to help”, and do not “approach things in a condescending manner.” However, she also identifies some men “who act like a know-it-all.” She also attempts to reason why they exhibit this behavior: “I don’t know if they do it because I’m a woman, but sometimes it feels like that, where they’re like mansplaining so often in and outside the classroom, there is definitely a culture of mansplaining.” For example, she recalls working with a male peer lab partner via Zoom. She says that her partner was extremely condescending and, “…I think he just thought he was smarter than everyone else in the room at all times and always had to be the one to answer the question first, always had to be the one to know it…I ended up switching my partners actually because I just could not deal with that.”
@@
While on Zoom, she says that her partner would dominate breakout room conversations and assignments, leaving her and her other woman partner behind: “He would always just solve it himself and then present the answer, without even giving us a chance to try and that really bothered me because it made me feel like…he thinks oh, because he knows it, that all of us cannot be smart enough to even try…and it made me feel like I was being talked down to from him, thinking that I was less smart than he was.” The fact that this happened over Zoom may have made this situation especially difficult to be in. Ruby also adds that her partner would not collaborate on the written report and would write it himself. She says, “he would act like I didn’t know anything that I was doing, when I was fully capable of doing it myself and that really made me feel…angry about it because… I love collaboration, I think it’s so important, and I really liked when I’m able to work with somebody, we’re able to combine our own knowledge and figure out new stuff, and with him, I felt that I was limiting my own knowledge, because I wasn’t getting access to even try. And I felt he was mansplaining things all the freaking time. I really did not like it.”
@@
When asked if she thought that this behavior was specifically towards women, Ruby says, “100% yes.” Again, she explains that she thought that the student thought that he was smarter than others, “but he also believed that he was helping in a way. He thought that, Oh, these women can’t do it themselves, I’ll help them, I’ll do it for them… But in my mind at least, I believe that’s why he did, that I think he thought of women in a sense, that they were less capable and more vulnerable than men, and so he wanted to be this savior.” Thus, in her experience, this male peer thought that he was helping women by answering for them and doing all the work, when in reality, it had the opposite effect.
@@
Cathy recalls her physics REU experiences at large research universities, some of them being very negative. In one of her REUs, she was the only woman in a group of nine students and she was not familiar with the research topic. As a result, she felt like she could not ask questions or reach out to others, thus making her feel excluded. She explains that: “everybody else knew so much more than me and they kind of talked down to me, sometimes it was just really overwhelming and intimidating and it definitely harmed my confidence a lot, because since I didn’t feel like I could ask questions, I really didn’t understand what was going on, and I got super behind, which made it so it was harder to ask questions, it’s just kind of the cycle and I don’t really feel like I learned anything in the end because I was just so overwhelmed the whole time.” Without any support from REU mentors, the fact that the other male students gave her the impression that they were more knowledgeable than her and they talked down to her made her intimidated to ask questions, potentially making her not benefit from the entire REU experience when a supporting environment would have made it very beneficial.
IV.2.3 Paulina as an island among her college peers
@@
She says that being a woman of color is complicated because, “people perceive me as being different in the classroom because I’m a person of color and also because I’m a woman.” She says these visible identities are apparent and “that’s…the first thing they see and it’s how they assess my value or what I can contribute to the classroom.” As one of the few people of color in her program, we note that this may be extremely difficult to navigate as she constantly thinks about how others are perceiving her. She also says that students simultaneously are judging her and assessing her intellect and the value of interacting with her and what she will bring to collaboration of physics problem solving. Similar issues have been documented in previous studies involving interviews with six physics graduate students who were Black women [21].
Paulina explains that she even feels judged amongst other women in her course, “there’s also been this expectation that I am just not up to par with them, because [of] my racial background and so I feel there’s some intersections in terms of my experiences in STEM with women…I definitely know that beyond the realm of just gender, race also plays a really big role in it.” In Paulina’s experience, she doesn’t necessarily find solidarity amongst other White women either because she is a woman of color. Since her university is both small and a PWI, finding support and solidarity in other students who look like her can be extremely difficult.
@@
Being ignored by peers: Much of Paulina’s experiences are centered around her working alone on assignments because she is ignored by her peers during group work. Below, we discuss how her intersecting marginalized identities negatively impact her throughout and play a central role in her interactions with peers. Paulina explains that her peers will not value or listen to her contribution in groups because they make assumptions based on her identity as a woman of color: “I think there’s this assumption that… because I am a woman of color in STEM, I don’t necessarily have the same background or knowledge as they [do], so a lot of the times I’ll bring [up] points and it’ll just be ignored and that’s something that I’ve had to deal with a lot.” From this account, it is clear that Paulina is aware of stereotypes her peers may believe in, which ultimately result in them not engaging with her appropriately during group work.
She reveals that this semester for the first time, she has been placed in a group with students who she has worked with before in her class. These students have similar physics backgrounds to Paulina and she feels that although there seems to be some effort to meet and work through assignments, it is difficult. She reflects upon working with these students outside of class that she has interacted with in class earlier, “[t]hey either stick to doing things on their own, or like they already have a background with other students in the class, so they’ll just reach out to them, and if we happen to be in the same group, they’ll usually just choose to staying engaged with the students they already know, rather than work together as a group with me who they might not be as familiar with.” Thus, according to Paulina, other students feel more comfortable with others who share their identity and whom they know more personally or have interacted with more. The fact that these students have all been placed in a group with Paulina to work outside of the class and she is still feeling excluded (both inside and outside of the classroom) is very concerning. This formation of a clique makes it difficult for Paulina to meaningfully collaborate with other peers (including other women) as they choose to talk to other students in other groups who they do know and have similar identities (e.g., gender and race), as opposed to her, who is placed in their group.
@@
Paulina explains that although some of her classes incorporate group work, there is no effort to communicate or work on problems outside of class. Even for the case mentioned earlier in which she has been placed in a group for studying outside the class, she feels that it does not work effectively, and she is left to work on her own. She says that as a result: “no one is reaching out to each other to talk and work things out together so by default I just feel like I have to end up doing things on my own.” Thus, even when she has been placed in a group to work outside of class, her own group does not communicate about when they should get together to study or do homework, so she resorts to working alone. It is unclear whether students who have been placed in Paulina’s group are meeting with other students with similar identity outside of the class to study and do homework. In particular, she is not aware of what is happening since nothing happens through the messaging system they were supposed to communicate through when they were placed in the same group for out of class study. Whether this lack of communication is directed towards Paulina because of her identity or caused by other reasons such as none of the students in the group she has been placed meeting regularly, it is interpreted as lack of community by her. Thus, she is unable to lean on her peers as resources and collaborators, and can only rely on herself.
@@
Inside the classroom, when she is ignored by her peers, she feels like she cannot rely on her group mates either. She feels like her only option is to reach out to her instructor and says she feels like, “I was expected to learn things on my own and if I didn’t verify [that my problem solving process was correct with the professor] it reflected badly on my own personal knowledge and understanding of the subject.” Throughout her interview, it is clear that Paulina is not only aware of how others negatively perceive her and exclude her, but it also appears to take a toll on her. For example, she feels like she has to overachieve and “…prove that I am just as knowledgeable as everyone else in the space.” Paulina’s experience is also consistent with other women of color doing more than just learning physics [22]. The fact that she feels that she has to overachieve and prove that she is knowledgeable like her peers is an additional burden on her because she does not want to confirm any stereotype others may have related to her identity [61].
@@
Because other students would not work with her even in class even though she was in a group with them, she could not complete in-class assignments during class like her peers did. She says that she would fall behind and be “left to my own devices” to work on problems. She also describes a cycle where she could not use her peers as a resource, i.e., ask them questions, and would have to finish the assignment outside of class. On the other hand, her peers would be able to complete assignments during class. Not only does she have to manage feelings of being excluded and negatively stereotyped by her peers, she also has to spend extra time outside of class to finish her class assignments. We can see how her multiple marginalized identities in physics can take up emotional and mental space that could be used for learning physics and problem solving. She has to use additional time and resources to complete her in-class assignments due to the exclusion by peers and navigate her physics journey alone.
@@
IV.3 Interactions with college instructors
In this section, we describe the experiences the women had with their college physics instructors. These experiences include Positive and Negative interactions and perceptions.
IV.3.1 Positive interactions and perceptions